1/1

N. N. 41,

The House of the Rising Sun

There is a house in New Orleans, They call the Rising Sun. It’s been the ruin of many poor girl, And me, o Lord, I’m one.

If I had listened to what my mother said, I’d been at home today. But I was young and foolish, o God, Let a rambler lead me astray.

Go tell my baby sister: Never do like I have done! But shun that house in New Orleans They call the Rising Sun!

I’m going back to New Orleans, My race is almost run. I’m going back to spend my life Beneath the Rising Sun.

Verkinto de tiu ĉi Angla poemo estas N. N. 41.

Arg-718-1442 (2012-12-09 22:40:30)

Pri tiu ĉi usona kanto vidu ankaŭ la vikipediejojn http: // de. wikipedia. org/ wiki/ The_House_of_ the_Rising_Sun kaj http: // en. wikipedia. org/ wiki/ The_House_of_the_Rising_Sun .

http://www.poezio.net/version?poem-id=718&version-id=1442